OpenStudy (anonymous):

6x+4y=42 -3x+3y=-6 elimination

OpenStudy (texaschic101):

6x + 4y = 42-- reduces to 3x + 2y = 21 now we have : 3x + 2y = 21 -3x + 3y = -6 -------------add 5y = 15 y = 15/5 y = 3 now sub 3 in for y in either of the equations to find x 3x + 2y = 21 3x + 2(3) = 21 3x + 6 = 21 3x = 21 - 6 3x = 15 x = 15/3 x = 5 check... -3x + 3y = -6 -3(5) + 3(3) = -6 -15 + 9 = -6 -6 = -6 (correct) SOLUTION IS : (5,3)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you so much